40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 22 01:45:41 2011, Keiko, Update, LSC, Locking status update PRCL-OLT.pdfMICH-OLT.pdf
    Reply  Thu Sep 22 04:49:14 2011, Anamaria, Update, LSC, Locking status update - Some Scripts, No Louck 
    Reply  Thu Sep 22 11:50:37 2011, Koji, Update, LSC, Locking status update 
Message ID: 5513     Entry time: Thu Sep 22 04:49:14 2011     In reply to: 5512
Author: Anamaria 
Type: Update 
Category: LSC 
Subject: Locking status update - Some Scripts, No Louck 

The scripts I wrote can be found in /users/anamaria/scripts/sensemat/

]There are two of them:

- one that sets all the switches, gains, frequencies, etc, then cycles through the various RFPDs I and Q into the LOCKIN signal, so as to see the sensing matrix.

- the second one is a matlab script that takes the crappy file tdsavg outputs and makes it into a cute mag/phase matrix.

They're quite primitive at this point, I've forgotten a lot of tcsh... may improve later. But could be useful later to someone else at least.

I don't think it's particularly the fault of the script that we can't measure the sensing matrix. We can slam on the excitation by hand, and it holds for a little while. I set a wait time for lock to adjust, and most times it just oscillates a bit for a few seconds. Also, the script turns on the excitation and it's done, the rest is just measurement, then turns it off at the end. So during the script, there's not much to deal with, except keeping the lowpass filters quiet when switching the signal to demod; but that doesn't go anywhere, so it definitely doesn't disturb the ifo. Turns out pressing the RSET clear history button needs a 2 to make it happen.

I think I might prefer to set the excitation to run, and then do the old retrieve-data-later-nds-matlab thing. I do not trust these measurements without coherence and a bit of variance study, given instabilities.

Point is... Even on carrier, the PRC lock is not stable by any means. Can barely turn on low freq boosts, every other lock. Until we fix the lock stability issue, there's not much to measure I guess.

Unfortunately, I don't know how to make that happen. Before we leave on Friday we could do a few sanity checks such as measuring the noise of the RFPDs vs ADC+whitening, which I may have said I would do; and perhaps setting up a couple OSAs, one on REFL, one on AS, to make sure we know what the sidebands are doing. Both of which Rana suggested at some point.

(There used to be a quote here from Keiko here but I got mad when it reformated my entire log to be one cluster- hence the look)

ELOG V3.1.3-